# Measuring the (Un)Faithfulness of Concept-Based Explanations

### v2

@@ B.5 ICE
The open-source implementation provided by the authors was used to evaluate this approach. ICE uses NMF to find CAVs. Due to the potentially large number of embeddings, we use the batched-version of NMF. Concept importance is found using the TCAV method proposed in [17]. Embeddings are projected to the concept space using sci-kit learn’s transform method. ICE was extended to multi-attribute prediction by reformulating the task as independent binary classification tasks. Thus, CAVs and concept importances were learned for each task independently.
@@ B.6 CRAFT
The open-source implementation provided in the Xplique toolbox [8] (by the same authors) was used to evaluate this approach. CRAFT uses NMF to find CAVs and introduces a Sobol-based sensitivity approach to find concept importance. Embeddings are projected to the concept space using the toolbox’s built-in transform method.
@@ B.7 ConceptSHAP
    Added more efficient concept importance scoring via Shapley Values. This is implemented using KernelSHAP ([23]) and is class-specific.
@@ B.8 MCD
Instead of representing concepts as single CAVs, MCD represents concepts as multi-dimensional subspaces. In [35], any concept $l$ for class $i$ is characterized by a representative basis $C^{i,l}=\{\mathbf{v}_{j}^{i,l}|j=1...d_{l}\}$, where $d_{l}$ denotes the dimensionality of the subspace. To ensure the union of all concepts span the entire feature space, they define the orthogonal complement concept $C^{i,\perp}=\text{span}(C^{i,1},...,C^{i,K})^{\perp}$, where $K$ is the number of concepts originally discovered, letting $C^{i,K+1}\equiv C^{i,\perp}$. Then, any embedding $\mathbf{h}$ can be uniquely decomposed onto the concept basis as:
@@
Since MCD is computationally expensive when trying to discover concepts on a large number of embeddings, we randomly sample 10,000 embeddings if the number of embeddings extracted for a given class is above 10,000. We use the open-source implementation provided by the authors.
C Adversarial Attack Details
We conducted a white-box adversarial attack using Projected Gradient Descent (PGD) [24] to evaluate the robustness of U-CBEMs under adversarial perturbations. The attack aims to maximize the EMD between the probability outputs of the model on the original image and the adversary.
A PGD attack with 100 iterations, step size $\alpha=1$, and perturbation bound $\epsilon=0.1$ was implemented to iteratively perturb input images while maximizing the EMD loss. The final perturbed image with the highest EMD loss (across all iterations) was used for evaluation. The original implementation of PGD assumes that the images are normalized from [0,1] to define the perturbation bound. Our images were standardized (following typical ImageNet processing), so we empirically selected an $\epsilon$ that resulted in minor perceptible changes to the reference image.

### v3

@@ B.5 ICE
The open-source implementation provided by the authors was used to evaluate this approach. ICE uses NMF to find CAVs. Due to the potentially large number of embeddings, we use the batched-version of NMF. Concept importance is found using the TCAV method proposed in [kimTCAV]. Embeddings are projected to the concept space using sci-kit learn’s transform method. ICE was extended to multi-attribute prediction by reformulating the task as independent binary classification tasks. Thus, CAVs and concept importances were learned for each task independently.
@@ B.6 CRAFT
The open-source implementation provided in the Xplique toolbox [fel2022xplique] (by the same authors) was used to evaluate this approach. CRAFT uses NMF to find CAVs and introduces a Sobol-based sensitivity approach to find concept importance. Embeddings are projected to the concept space using the toolbox’s built-in transform method.
@@ B.7 ConceptSHAP
    Added more efficient concept importance scoring via Shapley Values. This is implemented using KernelSHAP ([shapley]) and is class-specific.
@@ B.8 MCD
Instead of representing concepts as single CAVs, MCD represents concepts as multi-dimensional subspaces. In [vielhaben2023multidimensional], any concept $l$ for class $i$ is characterized by a representative basis $C^{i,l}=\{\mathbf{v}_{j}^{i,l}|j=1...d_{l}\}$, where $d_{l}$ denotes the dimensionality of the subspace. To ensure the union of all concepts span the entire feature space, they define the orthogonal complement concept $C^{i,\perp}=\text{span}(C^{i,1},...,C^{i,K})^{\perp}$, where $K$ is the number of concepts originally discovered, letting $C^{i,K+1}\equiv C^{i,\perp}$. Then, any embedding $\mathbf{h}$ can be uniquely decomposed onto the concept basis as:
@@
Since MCD is computationally expensive when trying to discover concepts on a large number of embeddings, we randomly sample 10,000 embeddings if there are greater than these many embeddings extracted for a given class. We use the open-source implementation provided by the authors.
B.9 HU-MCD
HU-MCD directly extends upon MCD, adopting the same framework of concept subspace, orthogonal complement, and unique decomposition onto the concept basis. The only material difference is in the concept discovery stage, where concepts are discovered on representations arising from image segments, instead of from the feature map. Therefore, we make the same adaptations as MCD for HU-MCD. We use the open-source implementation provided by the authors.